40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 26 17:31:17 2010, yuta, Update, CDS, fixed OPLEV stuff and MCL filters optlevMC2.png
    Reply  Tue Oct 26 21:27:02 2010, Jenne, Update, CDS, fixed OPLEV stuff and MCL filters 
Message ID: 3788     Entry time: Tue Oct 26 17:31:17 2010     Reply to this: 3789
Author: yuta 
Type: Update 
Category: CDS 
Subject: fixed OPLEV stuff and MCL filters 

(Joe, Yuta)

Background:

 We are currently working on getting rid of "white stuff" in MEDM screens.
 Today, we fixed OPLEV stuff, MCL filters, and time stamps.

What we did:
 1. Plugged in OPLEV cables to ADC2. (See this wiki page for wiring)

 2. Connected ADC2 and OPLEV in Simulink model and fixed MEDM screens for OPLEVs (Attached #1).

 3. Put MCL filters for BS,ITMX,ITMY,PRM,SRM.
  They don't need them, but just for getting rid of "white stuff."
  They are connected to the ground, so the outputs are always 0.

 4. Fixed "TIME_STRING"s in MEDM screens so that they show current time correctly.
  You only need to put text monitor with channel "C1:FEC-DCU_NODE_ID_TIME_STRING" into master files(DEFAULTNAME things) and run generate_master_screens.py.
  It will automatically sets DCU ID correctly!! (Great work, Joe!)
  See this wiki page for more info on making MEDM screens.

 5. Checked OPLEV for MC2 by pointing a laser pointer to QPD. (For MC2, OPLEV is just a transmission beam position monitor)
  Each quadrant looked like they are connected to the right channel numbers.

Plan:
 - figure out what C1:SUS-NAME_MODE_SW1 does and fix
 - fix Whitening, Dewhitening ON/OFF button in main MEDM screens, so that they switch every channels' filters
 - make a new screen for MC (like the old one C1IOO_ModeCleaner.adl)
 - create a new mark for new MEDM screens

Attachment 1: optlevMC2.png  28 kB  | Hide | Hide all
optlevMC2.png
ELOG V3.1.3-